# Gradient Guidance for Diffusion Models: An Optimization Perspective

### v1

@@ Appendix A Characterization for Output Distribution of Backward Process
Thus, we get $s_{\theta}(x_{t},t)$ is exactly the score of marginal distribution of $X_{t}$, i.e., $\nabla\log p_{t}(x_{t})=s_{\theta}(x_{t},t)$. According to the proof in Section B.1, we get $s_{\theta}(x_{t},t)+\texttt{G}_{loss}(x_{t},t)=\nabla\log p_{t}(x_{t}\mid y)$. Thus, the backward SDE turns out to be
@@
Since the initial distribution $p_{0}(x_{0}\mid y)$ of the forward process can also be obtained by (25) where we replace the initial distribution as $p_{T}(x_{T}\mid y)$. According to Girsanov theorem, we get the KL divergence between the terminal distribution $p_{T}^{\leftarrow}$ of (25) and $p_{0}(x_{0}\mid y)$.
$\text{KL}\left(p_{0}\mid p_{T}^{\leftarrow}\right)=-\mathbb{E}_{p_{0}}\left[\log\left(\frac{p_{T}}{\varphi}\right)\right],$
where $p_{0},p_{T}$ are short hands for $p_{0}(x_{0}\mid y)$ and $p_{T}(x_{T}\mid y)$, and $\varphi(\cdot)$ is the density for the standard normal distribution $\mathcal{N}\left(0,I_{D}\right)$. Since in the forward process, $p_{T}\to\varphi$ when $T\to\infty$, we have $\text{KL}\left(p_{0}\mid p_{T}^{\leftarrow}\right)\to 0$ when $T\to\infty$. We complete the first part of the lemma. As for the second part, if data reside in $\operatorname{Span}(A)$ following Assumption 1, we have $\bar{\mu}=A\bar{u}$ and $\bar{\Sigma}=A\bar{\Sigma}_{u}A^{\top}$, where $\bar{u}={\mathbb{E}}_{x\in\mathcal{D},u=A^{\top}x}[u]$, $\bar{\Sigma}_{u}={\mathbb{E}}_{x\in\mathcal{D},u=A^{\top}x}[\left(u-\bar{u}\right)\left(u-\bar{u}\right)^{\top}]$. Thus, the covariance matrix of $X^{\leftarrow}_{T}$ is
@@
Appendix B Omitted Proofs in Section 4
@@
2Contents   Throughout the proof, we omit the subscript $x_{t}$ of $\nabla$ in score function $\nabla\log p_{t}(x_{t})$ and conditional score function $\nabla\log p_{t}(x_{t}|y)$.
B.1 Proof of Lemma 1
@@ B.2 Proof of Theorem 1
It holds ${\mathbb{E}}[x_{0}|x_{t}]=\frac{1}{\alpha(t)}(x_{t}+h(t)\nabla\log p_{t}(x_{t}))$. Via the score decomposition under linear subspace data in Chen et al. (2023a, Lemma 1), we have
@@
Appendix C Omitted Proofs in Sections 5 and 6
@@
C.1 Proof of Theorem 2
@@
The proof is a special case of Theorem 3 in Appendix C.2, via setting the representation matrix $A=I_{D}$. ∎
C.2 Proof of Theorem 3
@@
C.3 Proof of Theorem 4
@@
where $h_{\lambda}(x):=f(x)-(\lambda/2)\left\lVert x-\bar{\mu}\right\lVert^{2}_{\bar{\Sigma}^{-1}}$. Define $x^{\star}_{A,\lambda}=\mathop{\mathrm{argmax}}_{x=Au}h_{\lambda}(x)$. With some similar steps in proof in Appendix C.2, by choosing $B_{k}\geq\zeta^{-4k}(1-\zeta)^{-2}$, together with ${z}_{0}=\bar{\mu}$, we get
@@
with $\eta=\displaystyle\frac{2}{L+2\lambda}$, $\zeta=\lvert 1-\eta(L+\lambda)\rvert$ and $M_{1}=2L\sqrt{(1+\eta^{2})\log\left(\displaystyle\frac{2K}{\delta}\right)}$. Also, we can get (30) as in proof in Appendix C.2. We restate it here:
@@ C.4 Auxiliary Lemma
See Bubeck et al. (2015, Lemma 3.11) for a proof. ∎

### v2

@@ Appendix B Characterization for Output Distribution of Backward Process
Thus, we get $s_{\theta}(x_{t},t)$ is exactly the score of marginal distribution of $X_{t}$, i.e., $\nabla\log p_{t}(x_{t})=s_{\theta}(x_{t},t)$. According to the proof in Section C.4, we get $s_{\theta}(x_{t},t)+\texttt{G}_{loss}(x_{t},t)=\nabla\log p_{t}(x_{t}\mid y)$. Thus, the backward SDE turns out to be
@@
The initial distribution $p_{0}(x_{0}\mid y)$ of the forward process can also be obtained by (18) where we replace the initial distribution as $p_{T}(x_{T}\mid y)$. According to the data processing inequality, we get the bound of the total variation distance between the terminal distribution $p_{T}^{\leftarrow}$ of (18) and $p_{0}(x_{0}\mid y)$:
$\text{TV}\left(p_{0},p_{T}^{\leftarrow}\right)\leq\text{TV}\left(p_{T},\varphi\right),$
where $p_{0},p_{T}$ are short hands for $p_{0}(x_{0}\mid y)$ and $p_{T}(x_{T}\mid y)$, and $\varphi(\cdot)$ is the density for the standard normal distribution $\mathcal{N}\left(0,I_{D}\right)$. Since in the forward process, $p_{T}\to\varphi$ when $T\to\infty$, we have $\text{TV}\left(p_{0},p_{T}^{\leftarrow}\right)\to 0$ when $T\to\infty$. We complete the first part of the lemma. As for the second part, if data reside in $\operatorname{Span}(A)$ following Assumption 1, we have $\bar{\mu}=A\bar{u}$ and $\bar{\Sigma}=A\bar{\Sigma}_{u}A^{\top}$, where $\bar{u}={\mathbb{E}}_{x\in\mathcal{D},u=A^{\top}x}[u]$, $\bar{\Sigma}_{u}={\mathbb{E}}_{x\in\mathcal{D},u=A^{\top}x}[\left(u-\bar{u}\right)\left(u-\bar{u}\right)^{\top}]$. Thus, the covariance matrix of $X^{\leftarrow}_{T}$ is
@@
Appendix C Additional Materials for Section 4
@@
2Contents
C.1 Score decomposition for subspace data
Under 1, the score function $\nabla\log p_{t}(x)$ decomposes to two orthogonal parts: an on-support component belonging to the subspace; and an orthogonal component. We recall this key result in Proposition 2, which later plays a key role in deriving subspace preserving guidance.
Proposition 2 (Score Decomposition for Subspace Data ([10] Lem. 1, Thm. 3)).
Under 1, the score function $\nabla\log p_{t}(x)$ decomposes as
$\nabla\log p_{t}(x)=\underbrace{A\nabla\log p_{t}^{\sf LD}(A^{\top}x)}_{\mathbf{s}_{\parallel}(A^{\top}x,t)\text{:~{}on-support~{}score}}-\underbrace{{h^{-1}(t)}\left(I_{D}-AA^{\top}\right)x}_{\mathbf{s}_{\perp}(x,t)\text{:~{}ortho.~{}score}}.$ (19)
C.2 Proof of Proposition 1
We give the proof of Proposition 1, which shows the failure of naive gradient guidance.
Proof.
Under 1, the score can be decomposed to terms parallel and orthogonal to $\operatorname{Span}(A)$ (Proposition 2). Applying naive guidance, we examine the orthogonal reverse process:
$\mathrm{d}X_{t,\perp}^{\leftarrow}=\left[\frac{1}{2}-\frac{1}{h(T-t)}\right]X_{t,\perp}^{\leftarrow}\mathrm{d}t+b(t)g\mathrm{d}t+\left(I_{D}-AA^{\top}\right)\mathrm{d}\overline{W}_{t}.$
Solving this SDE, we get the expectation of the final state following $\mathbb{E}[X_{T,\perp}^{\leftarrow}]=\int_{0}^{T}\exp\left(-\int_{0}^{t}h^{-1}(s)\mathrm{d}s\right)e^{t/2}b(T-t)g\mathrm{d}t$. For the schedule $h(t)=1-\exp(-\sqrt{t})$, we have the coefficient of direction $g$ is larger than $\int_{0}^{T}\exp(-t/2-2\sqrt{t})b(T-t)\mathrm{d}t>\int_{0}^{1}\exp(-5/2)b_{0}\mathrm{d}t>0$ where we can assume $T>1$. Thus, $\mathbb{E}[X_{T,\perp}^{\leftarrow}]\neq 0$. This means the generated sample is leaving the subspace, i.e., naive gradient guidance will violate the latent structure. ∎
C.3 Discussion for gradient-like guidance
To further clarify the derivation of our gradient guidance, we present a variant of Lemma 1.
Proposition 3.
Under Assumption 2, we have
where $\mathbb{E}[x_{0}|x_{t}]$ denotes the conditional expectation of $x_{0}$ given $x_{t}$ in the forward process (1), and $\beta(t)=\alpha(t)/(\sigma^{2}+g^{\top}\Sigma^{-1}\left(I_{D}+\alpha^{2}(t)/h(t)\cdot\Sigma\right)^{-1}g)$.
Remarks.
However, even for Gaussian distributions, as long as $\Sigma\neq I$, the term of (6) is no longer proportional to $g$ but becomes a pre-conditioned version of the gradient. We show the guidance above can maintain the subspace structure of data in the experiments Appendix E.
Under $q(t)\equiv 1$ in the forward process (1), we plot the theoretical choice of $\beta(t)$ and $\alpha(t),h(t)$ to $t$ in Figure 7. In practice, the choice of $\alpha(t),h(t)$ can vary and they are determined by the forward process used for pre-training; and $\beta(t)$ can be treated as a tuning parameter to adjust the strength of guidance.
The proof of Proposition 3 is included in the proof of Lemma 1.
C.4 Proof of Lemma 1
@@ C.5 Proof of Theorem 1
It holds ${\mathbb{E}}[x_{0}|x_{t}]=\frac{1}{\alpha(t)}(x_{t}+h(t)\nabla\log p_{t}(x_{t}))$. Via the score decomposition under linear subspace data in Chen et al. [10, Lemma 1], we have
@@
Appendix D Additional Materials for in Sections 5 and 6
@@
D.1 Convergence to Regularized Maxima in Mean
Here we present a more fundamental version of Theorem 2.
Theorem 4 (Convergence to Regularized Maxima in Mean).
@@ D.1 Convergence to Regularized Maxima in Mean
The proof is a special case of Theorem 2 in Appendix D.2, via setting the representation matrix $A=I_{D}$. ∎
D.2 Proof of Theorem 2
We first provide a proof sketch.
Proof Sketch
We observe that (24) resembles a gradient ascent update from $\mu_{k}\approx\bar{z}_{k}$ to $\mu_{k+1}$ corresponding to a regularzed optimization problem (23). In this regularized objective, the original objective $f(x)$ incorporates an additional proximal term with $\lambda:=1/\eta_{k}$. Therefore we can analyze the convergence of $\mu_{k}$ by following the classical argument for gradient optimization.
@@ D.2 Proof of Theorem 2
D.3 Proof of Theorem 3
We first provide a proof sketch.
Proof Sketch
The proof idea is similar to the proof of Theorem 2. For simplicity, we analyze the case where only the most recent sample batch $\mathcal{D}_{k}$ is merged with $\mathcal{D}_{0}$ for finetuning the score function. More specifically, we let $w_{k,i}=0$ for $0<i<k$ and $w_{k,0}=1-w_{k,k}$. Similar to the proof of Theorem 4, we obtain a recursive update rule given by
where $\bar{z}_{k}\approx\mu_{k}$ is the empirical mean of previous samples. This update rule also closely resembles the gradient ascent iteration for maximizing a regularized objective. A key difference here is that we can control the weights $w_{k,i}$ to reduce the impact of $\mathcal{D}_{0}$ and make the regularization term vanish to zero. Thus the mean $\mu_{k}$ eventually converges to the global maxima.
@@ D.3 Proof of Theorem 3
where $h_{\lambda}(x):=f(x)-(\lambda/2)\left\lVert x-\bar{\mu}\right\lVert^{2}_{\bar{\Sigma}^{-1}}$. Define $x^{\star}_{A,\lambda}=\mathop{\mathrm{argmax}}_{x=Au}h_{\lambda}(x)$. With some similar steps in proof in Appendix D.2, by choosing $B_{k}\geq\zeta^{-4k}(1-\zeta)^{-2}$, together with ${z}_{0}=\bar{\mu}$, we get
@@
with $\eta=\displaystyle\frac{2}{L+2\lambda}$, $\zeta=\lvert 1-\eta(L+\lambda)\rvert$ and $M_{1}=2L\sqrt{(1+\eta^{2})\log\left(\displaystyle\frac{2K}{\delta}\right)}$. Also, we can get (28) as in proof in Appendix D.2. We restate it here:
@@ D.4 Auxiliary Lemma
See Bubeck et al. [9, Lemma 3.11] for a proof. ∎
Appendix E Additional Materials for Experiments
[…]
E.1 Simulation
E.1.1 Experiment Setup
For linear data structure, we set the data’s ambient dimension as $D=64$ and the linear subspace dimension as $d=16$. The linear subspace is represented by an orthogonal matrix $A\in\mathbb{R}^{D\times d}$. We randomly generate a matrix $A$ and fix it once generated. After that, we sample a data point $X$ by first randomly sampling a latent variable $U\sim\mathcal{N}(0,I_{d})$ and computing $X=AU$. We independently sample a total of $65536$ data points as our pre-training data set. For nonlinear data structure, data are uniformly sampled from a unit ball in $\mathbb{R}^{64}$.
The objective functions considered in our experiments are $f_{1}(x)=10-(\theta^{\top}x-3)^{2}$ and $f_{2}(x)=5-0.5\lVert x-b\rVert.$ Here, $\theta$ and $b$ are randomly generated and fixed afterward. Since our data assumes a low-dimensional subspace representation, it is convenient to decompose $\theta$ into $\theta_{\perp}=(I-AA^{\top})\theta$ and $\theta_{\parallel}=AA^{\top}\theta$, representing the off-support and on-support components. We refer to $\frac{\lVert\theta_{\perp}\rVert}{\lVert\theta_{\parallel}\rVert}$ as the off/on-support ratio. Analogously, for a generated sample, we can also define its off/on-support ratio. Clearly, a small off/on-support ratio indicates close vicinity to the subspace.
Score Network Pre-training.
We utilize a version of the U-Net [51], with 14.8M trainable parameters. Note that this is a complicated network going beyond the linear score function class considered in our theories. Following the implementation of Denoising Diffusion Probabilistic Models (DDPM, Ho et al. [32]), we train the U-Net o estimate the score function $\nabla\log p_{t}$, via minimizing the score matching loss introduced in Eqn. (3). We discretize the backward process to have $200$ time steps as in Nichol and Dhariwal [44], and the U-Net is trained using our generated data set for 20 epochs. We use Adam as the optimizer, set the batch size as $32$, and set the learning rate to be $10^{-4}$. After the pre-training phase, we confirmed that the data subspace structure is well learned, as the generated samples using the pre-trained diffusion model have an average off/on-support ratio of $0.039$.
Implementation of Algorithm 1.
Implementation of Algorithm 2.
Algorithm 2 differs from Algorithm 1 in that it allows additional fine-tuning of the pre-trained score network. In practice, to update the score network incorporating newly generated data, one does not have to exactly solve (14) by re-training the full model from scratch. Instead, (14) can be viewed as a guideline that motivates more computationally efficient ways for updating the pre-trained score. It is a common practice to only fine-tune the weights of the old model by performing gradient descent over a few batches of newly generated data, which is similar to the spirit of (14). To be more specific, we adopt a computationally lightweight fine-tuning strategy: We only perform one Adam optimization step using the re-weighted loss given by Eqn. (14) with a batch of $32$ generated samples. We set the learning rate as $10^{-6}$. This simple strategy already demonstrates good performances as shown in Figure 10. Other implementation details are kept the same as those of Algorithm 1.
E.1.2 Results
We first demonstrate our gradient guidance $\texttt{G}_{loss}$ preserves the subspace structure learned from the pre-trained model. For comparison, we also tested the naive guidance G defined following Proposition 3 (with $\Sigma=I$). For a quick reference, we repeat the definition here:
Comparing G and $\texttt{G}_{loss}$ on Preserving Subspace Structure.
(a) Algorithm 1
(b) Algorithm 2
(c) 300-350 round of (a)
(d) 1000-1200 round of (b)
Algorithm 1 Converges to Regularized Optima.
(a) $\theta=A\beta^{*}$
(b) $\frac{\|{\theta_{\bot}}\|}{\|{\theta_{\|}}\|}=9$
(c) $b=4\cdot\textbf{1}_{D}$
(d) $b\sim\mathcal{N}(4\cdot\textbf{1}_{D},9\cdot I_{D})$
Algorithm 2 Converges to Global Optima.
Algorithm 2 converges to the maximal value of the objective function $f_{1}=10-(\theta^{\top}x-3)^{2}$ as shown in Figure 10(a). In Figure 10(b), we visualize the distribution of generated samples of Algorithm 1 (blue) and 2 (red), respectively, as the iteration evolves. We see that samples from Algorithm 1 mostly stay close to the pre-training data distribution (area described by the dotted contour). In constrast, samples of Algorithm 2 move outside the contour, as the diffusion model gets fine-tuned using self-generated data.
(a) Convergence of Algorithm 2
(b) Distribution of generated samples
Results for Nonlinear Data Structure.
We apply Algorithm 1 to data uniformly sampled from a unit ball in $\mathbb{R}^{64}$. The objective reward function is defined as $f(x)=\theta^{\top}x$, where $\left\lVert\theta\right\rVert=1$. The left panel of Figure 11 demonstrates that rewards increase and converge when using Algorithm 1. Higher guidance strength $\Delta$ (corresponding to lower regularization) results in a higher convergent reward. The right panel of Figure 11 shows that, for the same reward level, gradient guidance achieves a smaller deviation from the unit ball compared to the naive gradient approach. This suggests that gradient guidance can better preserve data structure for nonlinear manifolds.
Figure 11: Nonlinear data structure experiment. We apply Algorithm 1 to data uniformly sampled from a unit ball in $\mathbb{R}^{64}$. The objective reward function is $f(x)=\theta^{\top}x$, where $\left\lVert\theta\right\rVert=1$. Left: Rewards increase and converge with Algorithm 1. Higher guidance strength $\Delta$ (lower regularization) results in a higher convergent reward. Right: For the same reward, gradient guidance achieves a smaller deviation from the unit ball compared to the naive gradient. This indicates that gradient guidance can preserve data structure for nonlinear manifolds.
E.2 Image Generation
Hyperparameters.
Since tuning parameters $y$ and $\beta(t)$ are both for the strength of guidance, we can fix one of them. We choose $\beta(t)=100$ as suggested by [68], and set a series of $y\in\left\{2,4,10\right\}$. We run batch size $=20$ samples parallelly across iterations to evaluate the rewards. The prompt is uniformly sampled from the $1000$ classes of ImageNet [20].
E.3 Time Efficiency
We summarize the time cost of our experiments on one NVIDIA A100 GPU in Table 1.
Total runtime (iterations) Per iteration No guidance
Simulation 3.8 min (50 iter) 76 min (1000 iter) 4.6 s 2.6 s
Image 1.3 min (5 iter) 15.8 s 4.9 s
Table 1: Runtime Efficiency of Algorithm 1. Red refers to the total time to converge. No guidance refers to the time for one-time inference of the pre-trained model.